OpenStudy (anonymous):

Solve for v , where v is a real number. v-5=(square root -5v+31) (If there is more than one solution, separate them with commas.)

OpenStudy (dan815):

square both sides and simplify

OpenStudy (anonymous):

There aren't any solutions.

OpenStudy (dan815):

(v-5)^2=-5v+31 v^2-10V+25=-5v+31 v^2-5-6=0 (v-6)(v+1)=0 v=6,v=-1

OpenStudy (dan815):

There is a typo.

OpenStudy (dan815):

(v-5)^2=-5v+31 v^2-10V+25=-5v+31 v^2-5V-6=0 (v-6)(v+1)=0 v=6,v=-1

OpenStudy (anonymous):

Oh, derp. This is a radical equation right?

OpenStudy (goformit100):

A Warm Welcome to Opens Study

OpenStudy (anonymous):

ln(x+8)-ln 3=ln x solve for x.

OpenStudy (dan815):

ln rules, simplifies the left side try to figure it out :)

OpenStudy (dan815):

remember what this "ln" function really means